OpenStudy (anonymous):

Which value is an extraneous solution to the equation, |3x+2| = 4x + 5? -1 -3 -7 10

OpenStudy (calculusxy):

take out the absolute value bars for now. 3x+2=4x+5 now combine like terms. 4x-3x=1x => 2=1x+5 Solve for x. 2-5= -3 => -3=1x x=-3

OpenStudy (jdoe0001):

\(\bf |3x+2| = 4x + 5\implies \begin{cases} +(3x+2) = 4x + 5\implies 3x+2 = 4x + 5\\ \quad \\ \bf -(3x+2) = 4x + 5\implies -3x-2=4x+5 \end{cases}\) what values would you get for "x" from that?

OpenStudy (anonymous):

Thanks @calculusxy and @jdoe0001

OpenStudy (calculusxy):

Ur welcome :)